One number is 12 more than another number. If one-half of the smaller number equals one-third of the larger number, what is the larger number?
Answer by stanbon(75887)

One number is 12 more than another number. If one-half of the smaller number equals one-third of the larger number, what is the larger number?
----------------
x = y + 12
That means x is the larger number.
-----------------
(1/2)y = (1/3)x
------
Modify::
x = y + 12
x = (3/2)y
----------------------
y + 12 = (3/2)y
(1/2)y = 12
y = 24
x = 36
---------------
Cheers,
Stan H.
---------------